Power Drive Corporation designs and produces a line of golf equipment and golf

apparel. Power Drive has 100,000 shares of common stock outstanding and 5000

shares of preferred stock outstanding as of the beginning of 2018. Power Drive has the

following transactions affecting stockholders’ equity in 2018.

March 1Issues 65,000 additional shares of $1 par value common stock for $62 per share.

May 10 Repurchases 6,000 shares of treasury stock for $65 per share

May 21 Reissues 3,000 shares of treasury stock purchased on May 10 for $70 per share.

May 28 Issued 300 shares of preferred stock par value $50 8% for $ 85 per share.

June 1 Declares a cash dividend of $2.00 pershare to all stockholders of record on

June 15. (Hint:Dividends are not paid on treasury stock.)

July `Pays the cash dividend declared on June 1

1.Journalize the transactions listed above

REQUIRED PART 2

Answer the following questions

What is the dollar amount of the following accounts after the journal entries have been

made?

Preferred stock account

Common stock account

Treasury stock account

Cash dividend paid to the stock holders
Answer #1

Preferred stock account : 265000

Preferred stock
Balance as on 1Jan 2018 (5000*50) 250000
May 28 Issue (300*50) 15000
Ending Balance 265000

b)Common stock account :165000

Common stock
Balance as on 1Jan 2018 (100000*1) 100000
March 1 Issue (65000*1) 65000
Ending Balance 165000

c)Treasury stock account : 195000

Treasury stock
Balance as on 1 Jan 2018 0 May 21 Reissue (3000*65) 195000
May10 Repurchase (6000*65) 390000
Ending Balance 195000

d)cash dividend paid to stockholders:324000

Number of common shares outstanding = Beginning + additional issue on March 1 -Repurchase on May 10 -Reissue on May 21

        = 100000 +65000 -6000 +3000

        = 162000 shares

cash dividend paid = 162000*2 = $ 324000